# Particle migration in areas of constricted flow

### v1

@@ IV Conclusions
Acknowledgements.
This research has been supported by the Xunta de Galicia (Grant No. 2021-PG036) and Ministerio de Ciencia e Innovación (Grant No. PID2022-141626NB-I00).
AUTHOR DECLARATION
Conflict of interest
@@ Acknowledgements.
Author Contributions
Raquel Dapena-García: Methodology, Conceptualization, Validation, Formal analysis, Writing - Original Draft. Vicente Pérez-Muñuzuri: Conceptualization, Writing - Review &\&& Editing, Supervision, Funding acquisition, Project administration.
Data Availability Statement
The data that support the findings of this study are available from the corresponding author upon reasonable request.
@@ Appendix A Grid convergence details
$<Re_{p}>=\frac{1}{NT}\sum_{t,k}\frac{2R{\bf{v}}_{k}}{\nu}$ (15)
with ${\bf{v}}_{k}$ being the k-th particle velocity at time $t$. Particle radius ($R=5$, $10$, $15$, $20$), stenosis height $h_{0}$, and initial fluid velocity $u_{max}$ were renormalized appropriately as $H\times L$ was varied, and to keep the Reynolds number $Re=400$ constant with $\nu=0.04$, which was widely used in the results shown above. The number of time steps $T$ in each simulation was increased by a factor of 4 to simulate the same physical time in each case. $<Re_{p}>$ calculated for the three larger grids are within $\approx 4\%$ error as shown in Table 1, with the discrepancies being due to the random initial particles positions, and low number of particles. As mean particle velocities in the central flow increase with occlusion, $<Re_{p}>$ also does. Ultimately, in our simulations, to be computationally efficient, we have chosen the medium grid ($H\times L=400\times 4000$ cells).
$H\times L$ $<Re_{p}>(25\%)$ $<Re_{p}>(50\%)$

### v2

@@ 4 Conclusions
It should be noted that extending the present analysis to three dimensions could alter the quantitative outcomes. In such geometries, secondary flows, radial shear gradients, and out-of-plane particle motions can modify migration dynamics compared to the 2D case presented here. Blockage of blood flow in the constricted region due to particle clogging is unlikely in 3D vessels, reducing the likelihood of complete obstruction. In addition, particle orientation and contact area with the vessel wall are more complex in 3D, potentially influencing adhesion probabilities. While the qualitative trends reported here—such as the dependence of margination and adhesion on particle size and aspect ratio—are expected to remain valid, the precise migration rates and adhesion efficiencies may differ. A full three-dimensional study would therefore be an important direction for future work.
CRediT authorship contribution statement
Raquel Dapena-García: Methodology, Conceptualization, Validation, Formal analysis, Writing - Original Draft. Vicente Pérez-Muñuzuri: Conceptualization, Writing - Review &\& Editing, Supervision, Funding acquisition, Project administration.
Declaration of competing interest
@@
Data availability
Source code for the LBM simulations is available from the authors upon reasonable request.
Acknowledgements
This research has been supported by the Xunta de Galicia (Grant No. 2025-PG025) and Ministerio de Ciencia e Innovación (Grant No. PID2022-141626NB-I00).
@@ Appendix A Grid convergence details
$Re_{p}>=\frac{1}{NT}\sum_{t,k}\frac{2R{\bf{v}}_{k}}{\nu}$ (17)
with ${\bf{v}}_{k}$ being the k-th particle velocity at time $t$. Particle radius ($R=5$, $10$, $15$, 2020), stenosis height $h_{0}$, and initial fluid velocity $u_{max}$ were renormalized appropriately as $H\times L$ was varied, and to keep the Reynolds number $Re=400$ constant with $\nu=0.04$, which was widely used in the results shown above. The number of time steps $T$ in each simulation was increased by a factor of 4 to simulate the same physical time in each case. $Re_{p}>$ calculated for the three larger grids are within $\approx 4\%$ error as shown in Table 1, with the discrepancies being due to the random initial particles positions, and low number of particles. As mean particle velocities in the central flow increase with occlusion, $Re_{p}>$ also does. Ultimately, in our simulations, to be computationally efficient, we have chosen the medium grid ($H\times L=400\times 4000$ cells).
$H\times L$ $Re_{p}>(25\%) <$$Re_{p}>(50\%)$